OpenStudy (anonymous):

help!! The function ƒ(x) = x2 is not one-to-one. Find a portion of the domain where the function is one-to-one and find an inverse function.

OpenStudy (anonymous):

@CGGURUMANJUNATH hey you have helped me before! could you explain this to me?

OpenStudy (anonymous):

@iGreen any ideas?

OpenStudy (kaylaprincess):

Limit the domain to x ≥ 0. Inverse: standard procedure is to replace f(x) on the left by x, replace x by f^(-1) on the right, and solve for f^(-1): x = (f^(-1))² + a x - a = (f^(-1))² Recall that we have limited the domain of f (which is the range of f^(-1)) to non-negative argument; hence, we want the positive square root: √(x - a) = f^(-1) Source : https://answers.yahoo.com/question/index?qid=20130730072233AAzEAki

OpenStudy (kaylaprincess):

refresh page if comment shows up as symbols //

OpenStudy (anonymous):

thank you! but they arent actually the same problems. :(